Part of $20,000 is invested at 12% interest. The rest is invested at 8% interest. The total interest fro the two accounts is $1,800. How much money was invested at each rate?
Answer by checkley77(12844)

.12x+.08(20,000-x)=1,800
.12x+1600-.08x=1,800
.04x=1,800-1,600
.04x=200
x=200/.04
x=5,000 invested @ 12%.
20,000-5,000=15,000 invested @ 8%
Proof:
.12*5,000+.08*15,000=1,800
600+1,200=1,800
1,800=1,800